Disk VSN: HOB+1012 | Data volume at beginning: ? GB
- Issues with very large clkoff measurement, auto correlations, configuring dbbc. Have restarted / rebooted DBBC, mark5, hobart etc. Warren will look at it in the morning.
- 0537 - Experiment resumed after dealing with multiple issues. DBBC was in unusable state, apparently due to a fault in the 10 MHz reference cable. Also investigated RF/IF system following up on recentr non-detections. Found a failed cable compensator which has been bypassed. (JMc)
- Ongoing issue - No valid Tsys readings. Origin unclear. (JMc)
Disk VSN: HOB+0098 | Data volume at beginning: 10522 GB
- 1800UT Experiment started okay (Patrick).